jigglypuff314 (jigglypuff314):

The ratio of the circumference of a circle to twice the diameter of the circle is which of the following? A. pi/4 B. pi/2 C. pi D. 2pi E. 4pi I know that it's B... but idk how to do this. Help?

OpenStudy (anonymous):

circumference=2.pi.r , diameter =2r , r=radius 2.pi.r/2d=2.pi.r/4r=pi/2 C.pi/2 is correct

OpenStudy (math&ing001):

Circumference = pi * diameter So Circumference/(2*diameter) = pi/2

OpenStudy (moongazer):

The circumference is equal to pi times diameter. "pi*d" If you double the diameter, you will get "2*d" if you ratio it: pi*d/2*d you will get pi/2.

jigglypuff314 (jigglypuff314):

okay, got it! thanks guys! :D

OpenStudy (moongazer):

You're Welcome :)